OpenStudy (anonymous):

Solve: −4(x + 1) − 3 = −3(x − 4) −19 19 2 0

OpenStudy (anonymous):

-19

OpenStudy (anonymous):

-4(x+1)-3=-3(x-4) -4x-4-3=-3x+12 -4x-7=-3x+12 -1x-7=12 -1x=19 x=-19

OpenStudy (anonymous):

thanks guys,and girls :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

(: